the points (0,8/3), (1,3) and (82,30) are vertices  of which type of triangle ( acute ,obtuse,right, isoseles etc)

Grade:12

1 Answers

Pratham Ashish
14 years ago

the simplest way to determine the nature of the triangle is to compute the sides of it

 square of one side ,a^2 = { (0-1)^2 + (8/3 -3)^2  } = 10/9 =1.11

   b^2 = { (82-1)^2 + (30 -3)^2  } = 7290

 c^2 =  { (82-0)^2 + (30 -8/3)^2  } =7471.11

 it shows that , c^2 > a^2 + b^2

so the tiangle would be acute